OpenStudy (anonymous):

1+n squared/ n squared - 1 simplify

OpenStudy (asnaseer):

is this the question?\[\frac{(1+n)^2}{n^2-1}\]

OpenStudy (anonymous):

no its 1+ separate equation n squared over n square-1

OpenStudy (asnaseer):

so its this?\[\frac{1+n^2}{n^2-1}\]

OpenStudy (anonymous):

no the one is kinda one over one

OpenStudy (asnaseer):

or this?\[1+\frac{n^2}{n^2-1}\]

OpenStudy (valpey):

This?:\[1+\frac{n^2}{n^2-1}=\frac{n^2-1}{n^2-1}+\frac{n^2}{n^2-1}\]

OpenStudy (anonymous):

yes to asnaseer

OpenStudy (asnaseer):

ok, so you could simplify as follows:\[1+\frac{n^2}{n^2-1}=1+\frac{n^2-1+1}{n^2-1}=1+\frac{n^2-1}{n^2-1}+\frac{1}{n^2-1}=1+1+\frac{1}{n^2-1}\]

OpenStudy (anonymous):

thats confusing

OpenStudy (asnaseer):

which leads to:\[2+\frac{1}{n^2-1}\]

OpenStudy (asnaseer):

basically I added zero to the numerator by adding "-1+1"

OpenStudy (anonymous):

no valpey is right the way because you find the lcd which is nsquared minus one

OpenStudy (asnaseer):

it all depends on what you mean by /simplify/

OpenStudy (anonymous):

like find common denomenator and then you do it to the top to but then what i did was i got 2n squared minus one over nsquared minus one and i dont know what to do from there

OpenStudy (asnaseer):

I have given you my interpretation of simplify. I'm not sure how you have been taught to simplify by your teachers so I'm sorry but I cannot help out here.